# On anomalous dissipation induced by transport noise

### v2

@@ Appendix A Stochastic Meyers’ estimates
In this appendix, we prove maximal $L^{p}_{t}(L_{x}^{q})$-regularity estimates for parabolic SPDEs under only boundedness, measurability and parabolicity assumptions with $p$ and $q$ close to $2$. In the deterministic case, such estimates have been first proven by Meyers’ [64] via a perturbation argument. In the stochastic setting, such estimates have been proven in [15] where, due to the abstract setting, no extrapolation in space is given, i.e., $q=2$. Here, we give direct proof of the stochastic Meyers’ estimates following his original idea, which interestingly, also provides an improvement in the integrability in space. For simplicity, we mainly consider SPDEs on $\mathbb{T}^{d}$. However, extensions to domains and systems of parabolic SPDEs are possible.
Below, $(W^{n})_{n\geq 1}$ denotes a family of independent standard Brownian motions on a filtered probability space $(\Omega,\mathscr{A},(\mathscr{F}_{t})_{t\geq 0},\mathbf{P})$ and $\mathbf{E}\stackrel{{\scriptstyle{\rm def}}}{{=}}\int_{\Omega}\cdot\,\mathrm{d}\mathbf{P}$. Note that such Brownian motions might differ from the one introduced in Subsection 2.2.
@@ A.1. Parabolic stochastic Meyers’ estimates
(A.1) $\partial_{t}v&=\kappa\Delta v+f+\sum_{n\geq 1}\big{[}(\xi_{n}\cdot\nabla)v+g_{n}\big{]}\,\dot{W}_{t}^{n}&\text{ on }&\mathbb{T}^{d},\\ v(0,\cdot)&=0&\text{ on }&\mathbb{T}^{d},\end{aligned}\right.$
where $v$ is the unknown process and $f$, $\xi_{n}$ and $g_{n}$ are specified below. In the following, we say that a progressively measurable process $v:{\mathbb{R}}_{+}\times\Omega\to H^{1}(\mathbb{T}^{d})$ is a strong solution to (A.1) if a.s. $v\in L^{2}_{{\rm loc}}([0,\infty);H^{1}(\mathbb{T}^{d}))$, and for all $t\geq 0$,
@@
Here the equality is understood in $H^{-1}(\mathbb{T}^{d})$. Similar definitions are employed for the SPDEs considered in this section.
The following complements [8, Theorem 1.2] in case of $L^{\infty}$-transport noise.
@@
Let $\kappa>0$. Assume that $\xi_{n}=(\xi_{j}^{n})_{j=1}^{d}:{\mathbb{R}}_{+}\times\Omega\times\mathbb{T}^{d}\to{\mathbb{R}}^{d}$ is $\mathscr{P}\otimes\mathscr{B}(\mathbb{T}^{d})$-measurable for all $n\geq 1$, for which there exist $M_{0}>0$ and $\kappa_{0}\in[0,\kappa)$ such that, a.e. on +×${\mathbb{R}}_{+}\times\Omega\times\mathbb{T}^{d}$,
[…]
Then there exists $p_{0}(d,\kappa,\kappa_{0},M_{0})>2$ such that, for all $T\in(0,\infty)$, $p\in[2,p_{0}]$, $q\in[2,p]$ and progressively measurable processes $f$ and $g$ satisfying
$f\in L^{p}((0,T)\times\Omega;H^{-1,q}(\mathbb{T}^{d}))\quad\text{ and }\quad g=(g_{n})_{n\geq 1}\in L^{p}((0,T)\times\Omega;L^{q}(\mathbb{T}^{d};\ell^{2})),$
there exists a unique strong solution $v\in L^{p}((0,T)\times\Omega;H^{1,q}(\mathbb{T}^{d}))$ to (A.1) and
(A.4) $\|v\|_{L^{p}((0,T)\times\Omega;H^{1,q}(\mathbb{T}^{d}))}\lesssim_{d,\kappa,\kappa_{0},M_{0},T}$ $\,\|f\|_{L^{p}((0,T)\times\Omega;H^{-1,q}(\mathbb{T}^{d}))}$
$+$ $\,\|g\|_{L^{p}((0,T)\times\Omega;L^{q}(\mathbb{T}^{d};\ell^{2}))}.$
Some remarks are in order. Firstly, it is well-known that the condition (A.3) with $\kappa_{0}<\kappa$ is optimal in the parabolic regime. Secondly, by [4, Proposition 3.8], [68, Theorem 1.2] and the fact that the operator $v\mapsto-\Delta v$ on $H^{-1,q}(\mathbb{T}^{d})$ with domain $H^{1,q}(\mathbb{T}^{d})$ has a bounded $H^{\infty}$-calculus of angle 00 (due to the periodic version of [52, Theorem 10.2.25]), it follows that the LHS(A.4) can be replaced by
@@
The reader is referred to [52, Chapter 10] and [4, Subsection 2.2] for details on the $H^{\infty}$-calculus and Banach-valued fractional Sobolev spaces, respectively. The appearance of Besov spaces in (A.5) is optimal in light of the trace method, see e.g. [16, Section 3.12] and [3, Theorem 1.2]. Finally, by [4, Proposition 3.10], the estimates (A.4)-(A.6) also hold with non-trivial initial data $v(\cdot,0)\in L^{p}_{\mathscr{F}_{0}}(\Omega;B^{1-2/p}_{q,p}(\mathbb{T}^{d}))$ provided on the RHS(A.4) one also add $\|v(\cdot,0)\|_{L^{p}(\Omega;B^{1-2/p}_{q,p}(\mathbb{T}^{d}))}$.
@@
By a perturbation argument [8, Theorem 3.2], it is enough to prove the claim with $T=\infty$ and with $\Delta v$ replaced $\Delta v-v$. The advantage is that the latter operator is invertible. This allows us to work on the half-line ${\mathbb{R}}_{+}=(0,\infty)$ instead of intervals of finite length. In particular, this gives the independence of $p_{0}>2$ on $T$ of the maximal $L^{p}_{t}(L^{q}_{x})$-regularity estimates for (A.1). Hence, below, we consider
(A.7) $\partial_{t}v&=\kappa(\Delta v-v)+f+\sum_{n\geq 1}\big{(}\big{[}(\xi_{n}\cdot\nabla)v\big{]}+g_{n}\big{)}\,\dot{W}_{t}^{n}&\text{ on }&\mathbb{T}^{d},\\ v(0,\cdot)&=0&\text{ on }&\mathbb{T}^{d}.\end{aligned}\right.$
@@
Step 1: (Analysis of constants – case $\xi_{n}\equiv 0$ and $f\equiv 0$) For each $p>2$ there exists $K_{p}>0$ satisfying ${\lim_{p\downarrow 2}\sup_{2\leq r\leq p}K_{r}<1/\sqrt{2\kappa_{0}}}$ such that, for all $q\in[2,p]$ and progressively measurable process $g\in L^{p}({\mathbb{R}}_{+}\times\Omega;L^{q}(\ell^{2}))$,
(A.8) $\|v\|_{L^{p}({\mathbb{R}}_{+}\times\Omega;H^{1,q})}\leq K_{p}\|g\|_{L^{p}({\mathbb{R}}_{+}\times\Omega;L^{q}(\ell^{2}))},$
@@
where $v\in L^{p}({\mathbb{R}}_{+}\times\Omega;H^{1,q})$ is the unique strong solution to (A.7) with $\xi_{n}\equiv 0$ and $f\equiv 0$. The existence of a strong solution $v$ to (A.7) with $\xi_{n}\equiv 0$ for which (A.22) holds follows [69, Theorem 7.1] and the above mentioned boundedness of the $H^{\infty}$-calculus and invertibility of the operator $v\mapsto-\Delta v+v$ on $H^{-1,q}$ with domain $H^{1,q}$. For $p\geq 2$, let $K_{p}$ be the optimal constant constant for which (A.22) holds for all $q\in[2,p]$. Since $\kappa_{0}<\kappa$, to prove the claim of Step 1 it remains to prove that
(A.9) $\lim_{p\downarrow 2}\sup_{2\leq r\leq p}K_{r}\leq 1/\sqrt{2\kappa}$
To begin, note that $K_{2}\leq 1/\sqrt{2\kappa}$. The latter readily follows by an application of the Itô’s formula to $v\mapsto\|v\|_{L^{2}}^{2}$ (see e.g. [59, Theorem 4.2.5]) and taking expected values. Now, by complex interpolation (see, e.g., [51, Theorem 2.2.6]), for all $p>2$,
$K_{p_{\theta}}\leq K_{2}^{1-\theta}K_{p}^{\theta}\qquad\text{where}\qquad\tfrac{1}{p_{\theta}}=\tfrac{1-\theta}{2}+\tfrac{\theta}{p}.$
Hence, (A.9) follows from the above.
Step 2: Conclusion. To prove the claim of Theorem A.1, we employ the method of continuity [5, Proposition 3.13]. Hence, for $\lambda\in[0,1]$, consider the SPDE
@@
(A.10) $\partial_{t}v_{\lambda}&=\kappa(\Delta v_{\lambda}-v_{\lambda})+f+\sum_{n\geq 1}\big{(}\lambda\,\big{[}(\xi_{n}\cdot\nabla)v_{\lambda}\big{]}+g_{n}\big{)}\,\dot{W}_{t}^{n}&\text{ on }&\mathbb{T}^{d},\\ v_{\lambda}(0,\cdot)&=0&\text{ on }&\mathbb{T}^{d}.\end{aligned}\right.$
Note that, for $\lambda=1$, the above reduced to (A.7). Now, by [5, Proposition 3.13] it is enough to prove the existence of $p_{0}>2$ and $C_{0}>0$ such that, for all $p\in[2,p_{0}]$, $q\in[2,p]$, $\lambda\in[0,1]$ and progressively measurable processes $f\in L^{p}({\mathbb{R}}_{+}\times\Omega;H^{-1,q})$ and $g=(g_{n})_{n\geq 1}\in L^{p}({\mathbb{R}}_{+}\times\Omega;L^{q}(\ell^{2}))$,
(A.11) $\|v_{\lambda}\|_{L^{p}({\mathbb{R}}_{+}\times\Omega;H^{1,q})}\leq C_{0}(\|f\|_{L^{p}({\mathbb{R}}_{+}\times\Omega;H^{-1,q})}+\|g\|_{L^{p}({\mathbb{R}}_{+}\times\Omega;L^{q}(\ell^{2}))})$
where $v_{\lambda}\in L^{p}({\mathbb{R}}_{+}\times\Omega;H^{1,q})$ is a strong solution to (A.10). The key point is the independence of $C_{0},p_{0}$ on $\lambda\in[0,1]$.
@@
We begin with a reduction to the case $f\equiv 0$. Thus, let us assume that (A.11) holds with $f\equiv 0$. Since the operator $v\mapsto-\Delta v+v$ has a bounded $H^{\infty}$-calculus of angle 00, it also has deterministic maximal $L^{p}_{t}(L^{q}_{x})$-regularity by [72, Theorem 4.4.5]. Arguing as in [71, Theorem 3.9] (or by approximation), we can find a progressively measurable process $w\in L^{p}(\Omega;W^{1,p}({\mathbb{R}}_{+};H^{-1,q})\cap L^{p}({\mathbb{R}}_{+};H^{1,q}))$ which satisfies, a.s. for all $t\geq 0$,
$\partial_{t}w(t,\cdot)=\kappa(\Delta w(t,\cdot)-w(t,\cdot))+f(t,\cdot),\qquad w(0,\cdot)=0,$
@@
To prove (A.11) with $f\equiv 0$, let us note that (A.3) implies
$\|((\xi_{n}\cdot\nabla)v)_{n\geq 1}\|_{L^{q}(\ell^{2})}\leq\sqrt{2\kappa_{0}}\,\|\nabla v\|_{L^{q}}\ \ \text{ for all }\ v\in H^{1,q}.$
Hence, by Step 1, if $f\equiv 0$, then for all $\lambda\in[0,1]$, $p\in[2,\infty]$ and $q\in[2,p]$,
$\|v_{\lambda}\|_{L^{p}({\mathbb{R}}_{+}\times\Omega;H^{1,q})}$ $\leq K_{p}\big{(}\|g\|_{L^{p}({\mathbb{R}}_{+}\times\Omega;L^{q}(\ell^{2}))}+\|((\xi_{n}\cdot\nabla)v)_{n\geq 1}\|_{L^{p}({\mathbb{R}}_{+}\times\Omega;L^{q}(\ell^{2}))}\big{)}$
@@
Again, by Step 1, there exists $p_{0}(\kappa_{0},\kappa)>2$ such that $\sup_{2\leq p\leq p_{0}}K_{p}<1/\sqrt{2\kappa_{0}}$. Thus, the previous estimate yields, for all $2\leq p\leq p_{0}$ and $q\in[2,p]$,
@@
Hence, (A.11) with $f\equiv 0$ holds with a constant independent of $\lambda$ for $2\leq p\leq p_{0}$. ∎

### v3

@@ Appendix A Stochastic Meyers’ estimates
In this appendix, we prove maximal $L^{p}_{t}(L_{x}^{q})$-regularity estimates for parabolic SPDEs under only boundedness, measurability, and parabolicity assumptions with $p$ and $q$ close to $2$. In the deterministic case, such estimates have been first proven by Meyers’ [71] via a perturbation argument. In the stochastic setting, such estimates have been proven in [17] where, due to the abstract setting, no extrapolation in space is given, i.e., $q=2$. Here, we provide direct proof of the stochastic Meyers’ estimates following Meyers’ original idea, which, interestingly, also provides an improvement in the integrability in space. For simplicity, we mainly consider SPDEs on $\mathbb{T}^{d}$. However, extensions to domains and systems of parabolic SPDEs are possible.
Below, $(W^{n})_{n\geq 1}$ denotes a family of independent standard Brownian motions on a filtered probability space $(\Omega,\mathscr{A},(\mathscr{F}_{t})_{t\geq 0},\mathbf{P})$ and $\mathbf{E}\stackrel{{\scriptstyle{\rm def}}}{{=}}\int_{\Omega}\cdot\,\mathrm{d}\mathbf{P}$. We emphasize that these Brownian motions may differ from those introduced in Subsection 2.2.
@@ A.1. Parabolic stochastic Meyers’ estimates
(A.1) $\left\{\begin{aligned} \partial_{t}v&=\kappa\Delta v+f+\sum_{n\geq 1}\big[(\xi_{n}\cdot\nabla)v+g_{n}\big]\,\dot{W}_{t}^{n}&\text{ on }&\mathbb{T}^{d},\\ v(0,\cdot)&=0&\text{ on }&\mathbb{T}^{d},\end{aligned}\right$.
where $v$ is the unknown process and $f$, $\xi_{n}$ and $g_{n}$ are specified below. In the following, for a stopping time $\tau:\Omega\to[0,\infty]$, we say that a progressively measurable process $v:[0,\tau)\times\Omega\to H^{1}(\mathbb{T}^{d})$ is a strong solution to (A.1) (on $(0,\tau)$) if a.s. $v\in L^{2}_{{\rm loc}}([0,\tau);H^{1}(\mathbb{T}^{d}))\cap C([0,\tau);L^{2}(\mathbb{T}^{d}))$, and for all $t\in[0,\tau)$,
@@
Here, the above equality is understood in $H^{-1}(\mathbb{T}^{d})$. Similar definitions are applied to the other SPDEs considered in this section.
The following complements [9, Theorem 1.2] in case of $L^{\infty}$-transport noise.
@@
Let $\kappa>0$. Assume that
$\xi_{n}=(\xi_{j}^{n})_{j=1}^{d}:{\mathbb{R}}_{+}\times\Omega\times\mathbb{T}^{d}\to{\mathbb{R}}^{d}\text{ is $\mathscr{P}\otimes\mathscr{B}(\mathbb{T}^{d})$-measurable for all $n\geq 1$,}$
and that there exist $M_{0}>0$ and $\kappa_{0}\in[0,\kappa)$ such that, a.e. on ${\mathbb{R}}_{+}\times\Omega\times\mathbb{T}^{d}$,
(A.2) $\|(\xi_{n})_{n\geq 1}\|_{\ell^{2}}$ $\leq M_{0},$ $(\text{boundedness})$
(A.3) $\frac{1}{2}\sum_{n\geq 1}\big(\xi_{n}\cdot\eta\big)^{2}$ $\leq\kappa_{0}|\eta|^{2}\ \text{ for all }\eta\in{\mathbb{R}}^{d}.$ $(\text{parabolicity})$
Then there exists $p_{0}(d,\kappa,\kappa_{0},M_{0})>2$ such that, for all stopping times $\tau$ with values in $[0,T]$ with $T<\infty$, $p\in[2,p_{0}]$, $q\in[2,p]$ and progressively measurable processes $f$ and $g$ satisfying
$f\in L^{p}((0,\tau)\times\Omega;H^{-1,q}(\mathbb{T}^{d}))\quad\text{ and }\quad g=(g_{n})_{n\geq 1}\in L^{p}((0,\tau)\times\Omega;L^{q}(\mathbb{T}^{d};\ell^{2})),$
there exists a unique strong solution $v\in L^{p}((0,\tau)\times\Omega;H^{1,q}(\mathbb{T}^{d}))$ to (A.1) and
(A.4) $\|v\|_{L^{p}((0,\tau)\times\Omega;H^{1,q}(\mathbb{T}^{d}))}\lesssim_{d,\kappa,\kappa_{0},M_{0},T,p}$ $\,\|f\|_{L^{p}((0,\tau)\times\Omega;H^{-1,q}(\mathbb{T}^{d}))}$
$+$ $\,\|g\|_{L^{p}((0,\tau)\times\Omega;L^{q}(\mathbb{T}^{d};\ell^{2}))}.$
Some remarks are in order. Firstly, it is well-known that the condition (A.3) with $\kappa_{0}<\kappa$ is optimal in the parabolic regime. Secondly, by [5, Proposition 3.8], [75, Theorem 1.2] and the fact that the operator $v\mapsto-\Delta v$ on $H^{-1,q}(\mathbb{T}^{d})$ with domain $H^{1,q}(\mathbb{T}^{d})$ has a bounded $H^{\infty}$-calculus of angle 0 (due to the periodic version of [58, Theorem 10.2.25]), it follows that in case $\tau\equiv T$ the LHS(A.4) can be replaced by
@@
We emphasize that (A.5) also holds in the case $p=2$ (and hence $q=2$). In the latter case, (A.4)-(A.5) coincide with the usual $L^{2}$-estimate for parabolic SPDEs as $L^{2}(\mathbb{T}^{d})=B^{0}_{2,2}(\mathbb{T}^{d})$, see e.g., [65, Chapter 4].
The reader is referred to [58, Chapter 10] and [5, Subsection 2.2] for details on the $H^{\infty}$-calculus and Banach-valued fractional Sobolev spaces, respectively. The appearance of Besov spaces in (A.5) is optimal in light of the trace method, see e.g. [18, Section 3.12] and [4, Theorem 1.2]. Finally, by [5, Proposition 3.10], the estimates (A.4)-(A.6) also hold with non-trivial initial data $v(\cdot,0)\in L^{p}_{\mathscr{F}_{0}}(\Omega;B^{1-2/p}_{q,p}(\mathbb{T}^{d}))$ provided on the RHS(A.4) one also add $\|v(\cdot,0)\|_{L^{p}(\Omega;B^{1-2/p}_{q,p}(\mathbb{T}^{d}))}$.
@@
By a perturbation argument [9, Theorem 3.2], it is enough to prove the claim with $T=\infty$ and with $\Delta v$ replaced by $\Delta v-v$. The advantage is that the latter operator is invertible. This allows us to work on the half-line ${\mathbb{R}}_{+}=(0,\infty)$ instead of intervals of finite length. Moreover, as will be evident from the proof below, this also gives the independence of $p_{0}>2$ on $T$ in the maximal $L^{p}_{t}(L^{q}_{x})$-regularity estimate for (A.1). Hence, below, we consider
(A.7) $\left\{\begin{aligned} \partial_{t}v&=\kappa(\Delta v-v)+f+\sum_{n\geq 1}\big(\big[(\xi_{n}\cdot\nabla)v\big]+g_{n}\big)\,\dot{W}_{t}^{n}&\text{ on }&\mathbb{T}^{d},\\ v(0,\cdot)&=0&\text{ on }&\mathbb{T}^{d}.\end{aligned}\right$.
@@
Step 1: (Analysis of constants – case $\xi_{n}\equiv 0$ and $f\equiv 0$) For each $p>2$ there exists $K_{p}>0$ satisfying
(A.8) $\limsup_{p\downarrow 2}K_{p}<1/\sqrt{2\kappa_{0}}$
such that, for all $q\in[2,p]$, all stopping times $\tau:\Omega\to[0,\infty]$, and all progressively measurable process $g\in L^{p}((0,\tau)\times\Omega;L^{q}(\ell^{2}))$, there exists a unique strong solution $v$ to (A.7) with $\xi_{n}\equiv 0$ and $f\equiv 0$ such that $v\in L^{p}((0,\tau)\times\Omega;H^{1,q})$ and
(A.9) $\|v\|_{L^{p}((0,\tau)\times\Omega;H^{1,q})}\leq K_{p}\|g\|_{L^{p}((0,\tau)\times\Omega;L^{q}(\ell^{2}))}.$
@@
We emphasize that the case $\tau=\infty$ a.s. is allowed. The existence of a strong solution $v$ to (A.7) with $\xi_{n}\equiv 0$ for which (A.9) holds follows [76, Theorem 7.1] and the above mentioned boundedness of the $H^{\infty}$-calculus and invertibility of the operator $v\mapsto-\Delta v+v$ on $H^{-1,q}$ with domain $H^{1,q}$. For $p\geq 2$ and $q\in[2,p]$, let $N_{q,p}$ be the optimal constant constant for which (A.9) holds. By complex interpolation (see, e.g., [57, Theorem 2.2.6]), it follows that
(A.10) $N_{p,q}\leq N_{p,2}^{1-\eta}N_{p,p}^{\eta}\ \text{ where $\eta\in(0,1)$ satisfies }\ \tfrac{1}{q}=\tfrac{1-\eta}{2}+\tfrac{\theta}{p}.$
Recall that $\kappa_{0}<\kappa$. To prove the claim of Step 1, it remains to show that (A.8) with $K_{p}=\sup_{2\leq q\leq p}N_{p,q}$. To this end, as $\kappa_{0}<\kappa$, it is enough to show that
(A.11) $\lim_{p\downarrow 2}N_{p,2}=\lim_{p\downarrow 2}N_{p,p}=1/\sqrt{2\kappa}.$
We prove only that $\lim_{p\downarrow 2}N_{p,p}=1/\sqrt{2\kappa}$ as the other is similar. Again, by complex interpolation [57, Theorem 2.2.6], it follows that, for all $p_{0}>p$,
$N_{p,p}\leq N_{2,2}^{1-\theta}N_{p_{0},p_{0}}^{\theta}\ \text{ where $\theta\in(0,1)$ satisfies }\ \tfrac{1}{p}=\tfrac{1-\theta}{2}+\tfrac{\theta}{p_{0}}.$
We claim that
(A.12) $N_{2,2}\leq 1/\sqrt{2\kappa}$
It is clear that (A.12) yields second in (A.11) as $p\to 2$ implies $\theta\to 0$. Next, we prove (A.12). To this end, note that from Itô’s formula to $v\mapsto\|v\|_{L^{2}}^{2}$ (see e.g. [65, Theorem 4.2.5]) and an integration by parts, it follows that, a.s. for all $t>0$,
$\|v(t)\|_{L^{2}}^{2}+2\kappa\int_{0}^{t}\int_{\mathbb{T}^{d}}(|v|^{2}+|\nabla v|^{2})\,\mathrm{d}x\,\mathrm{d}t=\int_{0}^{t}\int_{\mathbb{T}^{d}}\|g\|_{\ell^{2}}^{2},\mathrm{d}x\,\mathrm{d}t+\mathcal{M}_{t},$
where $\mathcal{M}_{t}$ is an $L^{1}(\Omega)$-martingale starting at 0. Thus, taking expected values in the previous identity, discarding the quantity $\mathbf{E}\|v(t)\|_{L^{2}}^{2}$, and using that $\mathbf{E}\mathcal{M}_{t}=0$, we obtain the (quenched) energy inequality:
$2\kappa\,\mathbf{E}\int_{0}^{t}\int_{\mathbb{T}^{d}}(|v|^{2}+|\nabla v|^{2})\,\mathrm{d}x\,\mathrm{d}t\leq\mathbf{E}\int_{0}^{t}\int_{\mathbb{T}^{d}}\|g\|_{\ell^{2}}^{2}\,\mathrm{d}x\,\mathrm{d}t\ \ \text{ for all }t>0.$
Hence, (A.12) follows from the above by taking $t\to\infty$.
Step 2: Conclusion. To prove the claim of Theorem A.1, we employ the method of continuity [6, Proposition 3.13]. The method of continuity is a powerful tool in the theory of (S)PDEs (see, e.g., [9] and [86, p. 113–115]). In this context, one considers a family of problems that vary continuously with respect to a parameter $\lambda\in[0,1]$. The method allows one to deduce the existence and uniqueness of solutions at $\lambda=1$, provided these properties are known at $\lambda=0$, and a uniform a priori estimate for solutions holds for all $\lambda\in[0,1]$ (here, uniform means that the constants are independent of $\lambda$). In applications, the problem with $\lambda=0$ is typically simpler. In our case, it corresponds to equation (A.7) with $\xi_{n}\equiv 0$, as analyzed in Step 1. The original problem typically corresponds to $\lambda=1$. The key advantage of this approach is that, for $\lambda\in(0,1]$, one can assume a priori the existence of a solution to the $\lambda$-problem. In the context of linear (S)PDEs, such as those considered here, the existence and uniqueness of solutions can also be obtained a posteriori by leveraging the a priori estimate through a contraction principle (see [6, Proposition 3.13]).
Coming back to the proof, from the above discussion, for $\lambda\in[0,1]$, we consider the SPDE:
@@
(A.13) $\left\{\begin{aligned} \partial_{t}v_{\lambda}&=\kappa(\Delta v_{\lambda}-v_{\lambda})+f+\sum_{n\geq 1}\big(\lambda\,\big[(\xi_{n}\cdot\nabla)v_{\lambda}\big]+g_{n}\big)\,\dot{W}_{t}^{n}&\text{ on }&\mathbb{T}^{d},\\ v_{\lambda}(0,\cdot)&=0&\text{ on }&\mathbb{T}^{d}.\end{aligned}\right$.
Note that, for $\lambda=1$, the above reduces to (A.7), while for $\lambda=0$ it reduces to the one considered in Step 1.
We claim that there exist constants $p_{0}>2$ and $C_{0}>0$ such that, for all $p\in[2,p_{0}]$, $q\in[2,p]$, $\lambda\in[0,1]$, stopping time $\tau:\Omega\to[0,\infty]$, progressively measurable processes $f\in L^{p}((0,\tau)\times\Omega;H^{-1,q})$ and $g=(g_{n})_{n\geq 1}\in L^{p}((0,\tau)\times\Omega;L^{q}(\ell^{2}))$, and any strong solutions $v_{\lambda}$ to (A.13) in $L^{p}((0,\tau)\times\Omega;H^{1,q})$ and $\lambda\in[0,1]$, it holds that
(A.14) $\|v_{\lambda}\|_{L^{p}((0,\tau)\times\Omega;H^{1,q})}\leq C_{0}(\|f\|_{L^{p}((0,\tau)\times\Omega;H^{-1,q})}+\|g\|_{L^{p}((0,\tau)\times\Omega;L^{q}(\ell^{2}))}).$
The key point is the independence of $C_{0},p_{0}$ on $\lambda\in[0,1]$. As discussed above, the existence of such $v_{\lambda}$ is taken for granted, based on the method of continuity, which will be rigorously applied at the end of this step.
@@
Let $\tau$ be a fixed stopping time. To prove (A.14), we begin with a reduction to the case $f\equiv 0$. Therefore, let us assume that (A.14) holds with $f\equiv 0$. Since the operator $v\mapsto-\Delta v+v$ has a bounded $H^{\infty}$-calculus of angle 0, it also has deterministic maximal $L^{p}_{t}(L^{q}_{x})$-regularity by [79, Theorem 4.4.5]. Arguing as in [78, Theorem 3.9] (or by approximation), we can find a progressively measurable process $w\in L^{p}(\Omega;W^{1,p}(0,\tau;H^{-1,q})\cap L^{p}(0,\tau;H^{1,q}))$ which satisfies, a.s. and a.a. $t\in(0,\tau)$,
(A.15) $\partial_{t}w(t,\cdot)=\kappa(\Delta w(t,\cdot)-w(t,\cdot))+f(t,\cdot),\qquad w(0,\cdot)=0,$
@@
To prove (A.14) with $f\equiv 0$, note that, for all $v\in H^{1,q}$,
$\|((\xi_{n}\cdot\nabla)v)_{n\geq 1}\|_{L^{q}(\ell^{2})}=\Big(\int_{\mathbb{T}^{d}}\Big(\sum_{n\geq 1}|(\xi_{n}\cdot\nabla)v|^{2}\Big)^{q/2}\,\mathrm{d}x\Big)^{1/q}\leq\sqrt{2\kappa_{0}}\,\|\nabla v\|_{L^{q}},$
where the last inequality follows by applying (A.3) pointwise and with $\eta=\nabla v(x)$.
Hence, in case $f\equiv 0$, for all $\lambda\in[0,1]$, $p\in[2,\infty]$ and $q\in[2,p]$,
$\|v_{\lambda}\|_{L^{p}({\mathbb{R}}_{+}\times\Omega;H^{1,q})}$ $\leq K_{p}\big(\|g\|_{L^{p}({\mathbb{R}}_{+}\times\Omega;L^{q}(\ell^{2}))}+\|((\xi_{n}\cdot\nabla)v_{\lambda})_{n\geq 1}\|_{L^{p}({\mathbb{R}}_{+}\times\Omega;L^{q}(\ell^{2}))}\big)$
@@
where $K_{p}$ is the constant in (A.9) of Step 1. From the latter, there exists $p_{0}(\kappa_{0},\kappa)>2$ such that $\sup_{2\leq p\leq p_{0}}K_{p}<1/\sqrt{2\kappa_{0}}$. Thus, the previous estimate yields, for all $2\leq p\leq p_{0}$ and $q\in[2,p]$,
@@
Hence, (A.14) with $f\equiv 0$ holds with a constant independent of $\lambda$ for $2\leq p\leq p_{0}$.
Now, we are in a position to apply the method of continuity [6, Proposition 3.13]. Note that, from Step 1 and the reduction to the case $f\equiv 0$ as performed below (A.14), we have existence and uniqueness for the SPDE (A.13) with $\lambda=0$ for solutions in the class $L^{p}((0,\tau)\times\Omega;H^{1,q})$, where $\tau$ is any stopping time with values in $0,\infty]$. Thus, the conclusion of this step follows from the method of continuity in [6, Proposition 3.13] and the a priori estimate (A.14) that is uniform in $\lambda\in[0,1]$. ∎
Remark A.2.
Below are some comments on the proof of Theorem A.1.
  • The additional dissipative term $\kappa v$ in (A.7) is not essential for the proof but ensures that the constant $K_{p}$ in (A.9) remains independent of $T$. Without it, $K_{p}$ would depend on $T$, leading to a time-dependent $p_{0}$ in the second step and, consequently, to (A.14) holding only for $p\in[2,p_{0}]$, $q\in[2,p]$. Introducing $\kappa v$ and working on the half-line avoids this artificial dependence, as anticipated at the beginning of the proof of Theorem A.1.
  • Theorem A.1 also implies the existence of strong solutions to (A.7), and estimate (A.14) even holds for $\tau=\infty$. This is due to the term $\kappa v$, which—as previously noted—ensures that $p_{0}$ remains independent of the time interval. One can check that the estimate (A.1) holds if the processes $f$ and $g_{n}$ have mean zero, as it restores invertibility of the Laplacian on $\mathbb{T}^{d}$.